OpenStudy (anonymous):

What is the value of X in the following geometric sequence X , 64 , ___ , ___ , 8

OpenStudy (anonymous):

hi help me

OpenStudy (anonymous):

no

OpenStudy (anonymous):

not right

OpenStudy (anonymous):

it is 256

OpenStudy (anonymous):

but i don't know why

OpenStudy (anonymous):

128

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

you have the last term = 8 = ar^4 second term =64=ar divide these eq 1/2=r if ar=64 then a(first term)=64*2=128

OpenStudy (anonymous):

ohhh I see

OpenStudy (anonymous):

u got it now? or have more doubts

OpenStudy (anonymous):

yes I unddrstand now.Thanks